OpenStudy (anonymous):

Compute (f o g )(x), given the functions below: f(x) = (3x-4)/(4x-5) g(x) = 1/4x

OpenStudy (angela210793):

fog=(3*1/4x -4)*(4*1/4x -5) so u put g(x) instead of x in the F(x)

OpenStudy (anonymous):

did that; want to see how far is can factorize

OpenStudy (anonymous):

f(g(x))=\[(3/4x-4)/(4/4x-5)=(3-16x)/(4-20x)\]

OpenStudy (anonymous):

that isnt correct

OpenStudy (anonymous):

pull out a 4 in the denom

OpenStudy (anonymous):

thanks

OpenStudy (angela210793):

Is X in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

brackett that cannot be right

OpenStudy (angela210793):

(3-16x)/4x (3-16x) 4x 3-16x ------------=---------*-----------=---------- (4-20x)/4x 4x 4-20x 4-20x Brackett is right

OpenStudy (anonymous):

@akileez: there's x in the denominator of the fraction in both numerator and denomenator of the function, hence it'll get cancelled...hope its not confusing u :)

OpenStudy (anonymous):

one sec: redoing - thanks guys

OpenStudy (angela210793):

uw :)

OpenStudy (anonymous):

but, my answer choices dont reflect what you calc'd

OpenStudy (angela210793):

it doesn't? wht is it?

OpenStudy (anonymous):

a) (1/4) (4x-5)/(3x-4)

OpenStudy (anonymous):

b) (1/4)(12x^2-12x-5)/(4x-5)x

OpenStudy (anonymous):

c) (1/4)(16x-3)/(5x-1)

OpenStudy (anonymous):

d) ((4(3x-4)x))/(4x-5)

OpenStudy (anonymous):

e) (1/4)(16x+13)/(5x+4)

OpenStudy (angela210793):

it is C cause 4-20x=4(1-5x) and that is the only alternative tht has tht in denom.

OpenStudy (angela210793):

both up and down r divided with 4 so.... Wht do u think Brackett?

OpenStudy (anonymous):

that makes sense to me; i see where I misplaced the x in the denom before hand - thanks guys very helpful!

OpenStudy (angela210793):

^_^